OpenStudy (anonymous):

For which operations is the set {–1, 0, 1} closed? Choose all answers that are correct. A. addition B. division C. multiplication D. subtraction

OpenStudy (anonymous):

@BPDlkeme234

OpenStudy (anonymous):

@cwrw238

OpenStudy (michele_laino):

I think your answer its nt B, because \[\frac{ 1 }{ 0 }\] is not dfine, so cannotbelongto your set

OpenStudy (anonymous):

@paki

OpenStudy (michele_laino):

oops... is not defined, so it can not belong to your set

OpenStudy (anonymous):

but any thing times and divid by 0=0 right

OpenStudy (anonymous):

but you have to chose more than one

OpenStudy (michele_laino):

no, it is this: \[\frac{ 0 }{ any-thing }=0\]

OpenStudy (anonymous):

im confused?

OpenStudy (michele_laino):

for example: what is \[\frac{ 0 }{ 5 }=0:5=...\]

OpenStudy (anonymous):

0

OpenStudy (michele_laino):

so, 0:(any number)=0

OpenStudy (anonymous):

hi @pielover123, this moves you into a new dimension

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how @BPDlkeme234

OpenStudy (michele_laino):

now I think that your answer can not D, since -1-(1)=-2, and -2 doesn't belong to your set

OpenStudy (anonymous):

into the dimension of group theory

OpenStudy (anonymous):

ok but there is more then one answer

OpenStudy (michele_laino):

oops..can not be D...

OpenStudy (michele_laino):

there are answer A and C

OpenStudy (anonymous):

Which equations show that the set of whole numbers is closed under multiplication? Choose all answers that are correct. A. –1 • –1 = 1 B. 0 • –1 = 0 C. 0 • 2 = 0 D. 2 • 1 = 2

OpenStudy (michele_laino):

for exmple A can be one of your answers

OpenStudy (anonymous):

refresh if you see weird stuff

OpenStudy (anonymous):

ok so a d for the las answer then right

OpenStudy (anonymous):

*last

OpenStudy (anonymous):

Let G be a set and x a binary operation on G. Then (G, x) is a group if the following properties hold: G1 Closure: For all g, h element of G, g x h is an element of G

OpenStudy (anonymous):

what

OpenStudy (anonymous):

take the real numbers, if 2 numbers multiplied together still are parts of the real set then this is closure

OpenStudy (anonymous):

sorry two real numbers

OpenStudy (michele_laino):

I think C and D, because a whole number is a number great or equal to zero

OpenStudy (michele_laino):

oops...greater or equal to zero...

OpenStudy (anonymous):

the real question is do they remain in the same set of numbers (i.e closed)? since the original numbers are in the set of integers, that is what you need to ask! do the answers remain in the set of integers